OpenStudy (anonymous):

It takes 10 minutes to fill Alishas bathtub and 12 minutes to drain the water out. How long would it take to fill it with the drain accidentally left open?

OpenStudy (anonymous):

The sad thing is I know this is probably super obvious.

OpenStudy (aroub):

LOL! and the MORE sad thing is I dont know how to solve it! :P

OpenStudy (anonymous):

Hopefully someone comes in and enlightens us, haha.

OpenStudy (aroub):

yea.. I'm sure someone will come! :D

OpenStudy (anonymous):

fills at rate of 1/10 per minute, drains at rate of 1/12 per minute. combined rate is \[\frac{1}{10}-\frac{1}{12}=\frac{1}{60}\] per minute

OpenStudy (anonymous):

so it will take 60 minutes = 1 hour to fill

OpenStudy (anonymous):

Wow yeah that was really simple. Thanks for showing me how to set that up my friend!